# Decoupled Classifier-Free Guidance for Counterfactual Diffusion Models

### v2

@@ 3.2 Decoupled Classifier-Free Guidance
The final guided score used in the diffusion model is computed as:
@@
The proposed DCFG framework is highly flexible, as it allows arbitrary groupings of attributes, regardless of whether attributes within a group are mutually independent or not. The only assumption required is that different groups are conditionally independent given the latent variable $\mathbf{x}_{t}$. This flexibility enables a wide range of configurations. For instance, setting $M=1$ recovers standard global CFG, while increasing $M$ provides progressively finer-grained control, including per-attribute guidance ($M=K$) as an extreme case where we assume all attributes are independent of each other.
3.3 DCFG for Counterfactual Generation
@@
We now detail how DCFG is integrated into DDIM-based counterfactual inference.
@@
Abduction. The abduction step proceeds as in eq. 4, where the conditioning vector $\mathbf{c}$ is obtained by embedding the semantic parent attributes $\mathbf{pa}$ using the attribute-split encoder defined in eq. 12.
@@
Prediction. The prediction step uses the DCFG-guided reverse DDIM trajectory:
$\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-{\alpha}_{t}}\cdot\boldsymbol{\epsilon}_{\mathrm{DCFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}})\right),$ (18)
$\mathbf{x}_{t-1}=\sqrt{{\alpha}_{t-1}}\cdot\hat{\mathbf{x}}_{0}+\sqrt{1-{\alpha}_{t-1}}\cdot\boldsymbol{\epsilon}_{\mathrm{DCFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}}),$ (19)
where $\boldsymbol{\epsilon}_{\mathrm{DCFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}})$ is computed as in eq. 16 using counterfactual conditioning embedding $\tilde{\mathbf{c}}$.
For the counterfactual generation task, we follow a two-group partitioning of attributes based on the causal graph. The affected group $\mathbf{pa}^{\text{aff}}$ consists of attributes that are directly intervened on, along with their causal descendants. The invariant group $\mathbf{pa}^{\text{inv}}$ comprises attributes that are not affected by the intervention and are expected to remain unchanged. These groups are assumed to be conditionally independent given the latent $\mathbf{x}_{t}$, consistent with the d-separation implied by the post-intervention causal graph. Under this setup, eq. 16 is instantiated with $M=2$ groups, applying separate guidance weights $\omega_{\text{aff}}$ and $\omega_{\text{inv}}$ to the affected and invariant groups, respectively. In practice, we find that $\omega_{\text{inv}}=1.0$ is typically sufficient when $\omega_{\text{aff}}$ is small. However, as $\omega_{\text{aff}}$ increases, the resulting intervention may introduce unintended drift in $\mathbf{pa}^{\text{inv}}$ due to more aggressive changes to the image. In such cases, raising $\omega_{\text{inv}}$ can help maintain the desired invariance. Determining the optimal $(\omega_{\text{aff}},\omega_{\text{inv}})$ pair remains an open problem for future exploration. Fig. A.8 visualizes the effect of $\omega_{\text{inv}}$.
@@ 4 Experiments
In this section, we demonstrate the benefits of the proposed approach across three public datasets. For each dataset, we train a diffusion model with the same architecture and training protocol, detailed in appendix C. We compare our DCFG against the standard CFG baseline. In all results, settings labeled as $\omega=X$ correspond to standard classifier-free guidance (CFG) with a global guidance weight. In contrast, configurations denoted by $\omega_{\text{aff}}=X$, $\omega_{\text{inv}}=Y$ represent our proposed DCFG, where separate guidance weights are applied to the intervened and invariant attribute groups, respectively. Following Melistas et al. [30], Monteiro et al. [32], we evaluate counterfactual quality using two metrics. Effectiveness ($\Delta$): Measured by a pretrained classifier as the change in AUROC for intervened attributes relative to $\omega=1.0$ (no CFG). Higher $\Delta$ indicates stronger intervention effect; large $\Delta$ on invariant attributes indicates unintended amplification. Reversibility: Assesses how well counterfactuals can be reversed to the original image using inverse interventions. We report MAE and LPIPS; lower values indicate better identity preservation. See section A.2 for details.
4.1 CelebA
Figure 1: Comparison of $\Delta$ metrics under different interventions in CelebA-HQ. Left: Intervention on Smiling. Right: Intervention on Young. Both use baseline $\omega=1.0$. Under global CFG, increasing $\omega$ boosts the intended attribute but amplifies non-target attributes. DCFG achieves similar improvements on the target attribute while mitigating amplification. See section D.2 for full quantitative results.
We begin our evaluation on the CelebA-HQ dataset [20], using Smiling, Male, and Young as independent binary parent attributes. Although some variables (e.g., Smiling and Mouth Open) may be causally linked in the real world, we deliberately adopt this simplified setup to isolate and diagnose inference-time failures caused by global CFG. Under this assumption, any unintended change in non-intervened attributes can be confidently attributed to attribute amplification by standard CFG, rather than valid downstream causal effects. Refer to section D.1 for more details.
Fig. 1 presents the $\Delta$ metrics under different guidance strategies for two separate interventions: do(Smiling) and do(Young). As the global guidance weight $\omega$ increases (left to right side of each plot), the $\Delta$ of the intervened attribute improves, but so do the $\Delta$ values of attributes that should remain invariant, indicating undesirable amplification. In contrast, the right side of each plot shows results for DCFG, where distinct weights are applied to affected ($\omega_{\text{aff}}$) and invariant ($\omega_{\text{inv}}$) attribute groups. This decoupled formulation achieves comparable or stronger improvement on the intervened attribute while keeping the others stable, validating DCFG ’s ability to produce more disentangled and effective counterfactuals.
Figure 2: Counterfactual generations in CelebA-HQ ($64\times 64$). Each row compares global CFG (left) and DCFG (right) across guidance weights. Top: global CFG causes amplification of Smiling under do(Male); Middle: do(Young) suppresses Male (i.e. amplifies $\texttt{Male}=no$); Bottom: do(Smiling) makes the subject appear older, adds glasses, and alters identity. DCFG mitigates these unintended changes and preserves non-intervened attributes. Refer to section D.3 for more visual results.
Fig. 2 illustrates how global CFG can introduce unitended changes by uniformly amplifying all conditioning signals, even when only one attribute is meant to change. In the top row, applying do(Male=no) with increasing $\omega$ inadvertently amplifies Smiling; in the middle row, do(Young=no) reduces Male expression; and in the bottom row, do(Smiling=yes) introduces changes to age, identity, and even adds glasses. These unintended shifts stem from global CFG treating all attributes equally. In contrast, DCFG applies decoupled guidance across attributes—assigning stronger weights to those affected by the intervention, allowing attributes that were not targeted by the intervention to remain unchanged. This results in counterfactuals that more faithfully reflect the intended change while preserving identity and consistency in non-intervened factors.
@@ 4.1 CelebA
Figure 3: Reversibility analysis in CelebA-HQ $(64\times 64)$. Left: Quantitative evaluation of how well the original image is recovered after generating a counterfactual and mapping it back to the original condition under do(Smiling). Right: A qualitative example showing a counterfactual generated under do(Male) and its reconstruction after reversing the intervention with CFG and our DCFG.
Fig. 3 evaluates the reversibility of counterfactuals in CelebA-HQ. The left panel reports quantitative metrics (MAE and LPIPS) that assess how well the original image is recovered after applying an intervention (e.g., do(Smiling)) and then reversing it. With global CFG, reversibility consistently degrades as guidance strength increases, with higher error values across both metrics. In comparison, DCFG yields lower MAE and LPIPS for the same guidance levels on intervened variables (i.e., comparing $\omega$ with $\omega_{\text{aff}}$), demonstrating improved ability to recover the original image. The right panel presents a qualitative example where a counterfactual is generated under do(Male) and then reversed. Notably, the reversed image appears older than the original—an artifact not of the intervention itself, but of global CFG’s tendency to amplify all attributes in the counterfactual parent vector. While the model reverts the intervened attribute, non-intervened ones like Young = no remain amplified, leading to unintended and compounding changes. In contrast, the proposed DCFG distinguishes between intervened and invariant attributes, applying stronger guidance only to the target group. This mitigates attribute amplification in non-target attributes during both generation and reversal, enabling more faithful, disentangled, and reversible counterfactuals.
4.2 EMBED
We evaluate DCFG on the EMBED [18] breast mammography dataset. We define a binary circle attribute based on the presence of circular skin markers, and a binary breast density label, where categories A and B are grouped as low and categories C and D as high. See section E.1 for details.
Fig. 4 presents results for counterfactual generation on EMBED. The bar plot on the left reports $\Delta$ effectiveness metrics, measuring how classifier performance changes relative to the baseline. While global CFG improves effectiveness for the target attribute (circle), it also increases effectiveness on non-intervened attributes such as density, indicating unintended attribute amplfication. DCFG mitigates this by applying selective guidance, maintaining stable performance on non-target attributes. The figure on the right illustrates a key example: applying do(density) under global CFG unintentionally amplifies the presence of circular skin markers, as evidenced by the increased number of visible circles in both the counterfactual and reversed images. This is suppressed under DCFG, where circle features remain unchanged in both counterfactual and reversed images.
@@ 4.2 EMBED
Figure 4: Evaluation of counterfactual generation on EMBED ($192\times 192$). Left: $\Delta$ metrics showing the effect of do(circle). DCFG improves target intervention effectiveness while suppressing spurious shifts in non-intervened attributes. Right: A visual example showing the input image, the counterfactual under do(density), the reversed image, and their difference maps (CF/Rev. - input). See section E.2 for full quantitative results and section E.3 for more visual results.
4.3 MIMIC
We evaluate our method on the MIMIC-CXR dataset [19]. We follow the dataset splits and filtering protocols from [9, 13], and focus on the binary disease label of pleural effusion. The underlying causal graph in De Sousa Ribeiro et al. [9] includes four attributes: race, sex, finding, and age. We adopt this setup, but since our goal is to study attribute amplification caused by CFG, we focus on sex, race, and finding, which we assume to be mutually independent for analysis purposes. See section F.1 for details.
Fig. 5 presents an evaluation of counterfactual generation in MIMIC-CXR, highlighting the advantages of our proposed DCFG. The bar plot on the left shows $\Delta$ metrics that quantify the change in effectiveness relative to the baseline $\omega=1.0$. While global CFG improves effectiveness for the intervened variable (finding), it also introduces substantial shifts in non-intervened attributes such as race and sex, revealing unwanted attribute amplfication. In contrast, DCFG achieves comparable or higher target effectiveness while suppressing spurious changes, demonstrating more precise and controlled generation. On the right, we show a qualitative example of a counterfactual generated under do(finding), its reversed reconstruction, and their corresponding difference maps. Compared to global CFG, our method yields localized, clinically meaningful changes in counterfactuals and better identity preservation in the reversed image—further supporting its robustness and causal faithfulness.
@@ 4.3 MIMIC
Figure 5: Evaluation of counterfactual generation on MIMIC ($192\times 192$). Left: $\Delta$ metrics showing the effect of do(finding). DCFG improves target intervention effectiveness while suppressing spurious shifts in non-intervened attributes. Right: A visual example showing the input image, the counterfactual under do(density), the reversed image, and their difference maps (CF/Rev. - input). See section F.2 for full quantitative results and section F.3 for more qualitative results.

### v5

@@ 2.1 Diffusion Models for Counterfactual Inference
where $t{=}T,\dots,1$, $\hat{\mathbf{x}}_{0}$ is the predicted clean image, and the final output $\tilde{\mathbf{x}}_{0}$ is the predicted counterfactual $\tilde{\mathbf{x}}$. In practice, decoding under the counterfactual condition $\tilde{\mathbf{c}}$ using the conditional denoiser alone may be insufficient for producing effective counterfactuals. Additional guidance is often required to steer generation toward the desired intervention (Sanchez and Tsaftaris, 2022; Sanchez et al., 2022a; Komanduri et al., 2024; Fontanella et al., 2024; Weng et al., 2024; Song et al., 2024; Pérez-García et al., 2024; Rasal et al., 2025; Kumar et al., 2025). In line with previous work, we adopt classifier-free guidance to improve intervention effectiveness, by which we mean the extent to which the generated counterfactual reflects the intended target-attribute change, rather than a causal treatment effect.
2.2 Classifier-Free Guidance
Classifier-free guidance (CFG) (Ho and Salimans, 2022) is a widely adopted technique in conditional diffusion models. It enables conditional generation without requiring an external classifier by training a single denoising model to operate in both conditional and unconditional modes. During training, the same denoising network is exposed to both conditional and unconditional inputs by randomly replacing $\mathbf{c}$ with a null token $\varnothing$, thereby learning conditional and unconditional score estimates. At inference time, CFG biases the sampling process toward regions more consistent with the conditioning signal, which can be understood as sampling from a reweighted conditional distribution:
$p^{\omega}(\mathbf{x}_{t}{\mid}\mathbf{c})\propto p(\mathbf{x}_{t})p(\mathbf{c}{\mid}\mathbf{x}_{t})^{\omega},$ (6)
where $\omega{\geq}0$ controls the guidance strength. This corresponds to interpolating between the unconditional and conditional scores:
$\nabla_{\mathbf{x}_{t}}\log p^{\omega}(\mathbf{x}_{t}{\mid}\mathbf{c})=\omega\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}{\mid}\mathbf{c})+(1{-}\omega)\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}).$ (7)
In practice, this is implemented by combining the model’s predictions with and without conditioning:
$\boldsymbol{\epsilon}_{\text{CFG}}(\mathbf{x}_{t},t,\mathbf{c})=(1{-}\omega){\cdot}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\varnothing)+\omega{\cdot}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c}).$ (8)
@@ 2.2 Classifier-Free Guidance
$\mathbf{x}_{t-1}$ $)\displaystyle{=}\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\Big(\mathbf{x}_{t}{-}\sqrt{1{-}\alpha_{t}}\,\boldsymbol{\epsilon}_{\text{CFG}}(\mathbf{x}_{t},t,\tilde{\mathbf{c}}$$\Big)$ (9)
[…]
Despite its effectiveness, CFG applies a single global guidance weight $\omega$ uniformly across the entire counterfactual embedding $\tilde{\mathbf{c}}$, which typically encodes multiple semantic attributes. In counterfactual generation, however, only the attributes affected by the intervention should be strengthened, while attributes outside the causal effect of the intervention should remain stable. Applying the same guidance strength to all components of $\tilde{\mathbf{c}}$ therefore couples intervention-relevant and intervention-invariant attributes during sampling. This can induce unintended changes to invariant attributes, a failure mode known as attribute amplification (Xia et al., 2024). Such off-target changes violate the intended relationships in the associated causal graph and undermine the axiomatic soundness of inferred counterfactuals (Monteiro et al., 2023).
To address this limitation, we propose a structured alternative that factors the guidance signal across semantically or causally defined attribute groups, allowing each group to be controlled with its own guidance weight.
3 Factored Classifier-Free Guidance
In this section, we present our Factored Classifier-Free Guidance (FCFG) for counterfactual image generation. We first propose a simple attribute-split conditioning embedder (Sec. 3.1) as a practical implementation that separates attributes in the embedding space to enable selective control. Building on this, we then describe our FCFG formulation in detail (Sec. 3.2), which allows distinct guidance strengths to be applied to different subsets of attributes within an assumed causal graph. Finally, we present how FCFG is integrated into DDIM-based counterfactual inference (Sec. 3.3), detailing its application across abduction, action, and prediction steps using causally defined attribute groupings.
3.1 Attribute-Split Conditioning Embedding
In practice, raw conditioning inputs such as discrete image labels or structured attributes (e.g., a patient’s sex, race, or disease status) are not used directly in diffusion models but transformed into dense vectors using embedding functions, typically via multi-layer perceptrons (MLPs) (Dhariwal and Nichol, 2021), convolutional encoders (Zhang et al., 2023), or transformer-based text encoders (Ho and Salimans, 2022; Ramesh et al., 2022). These embeddings align semantic or categorical inputs with the model’s internal representations, but conventional designs often entangle multiple attributes into a single conditioning vector, making it difficult to independently control attributes during sampling.
To address this, we introduce a simple attribute-split conditioning embedding technique that preserves the identity of each attribute in the embedding space. Let $pa_{i}$ denote the raw value of the $i$-th parent attribute (e.g., a binary indicator or scalar). Each $pa_{i}$ is embedded independently via a dedicated MLP: $\mathcal{E}_{i}:\mathbb{R}^{d_{i}}{\rightarrow}\mathbb{R}^{d}$, and the final condition vector is formed by concatenating the outputs:
$\mathbf{c}$ $=\operatorname{concat}\!\big(\mathcal{E}_{1}(pa_{1}),\mathcal{E}_{2}(pa_{2}),\dots,\mathcal{E}_{K}(pa_{K})\big),$ (10)
where $\mathbf{c}{\in}\mathbb{R}^{Kd}$. This architecture provides a flexible representation where each attribute occupies a separate block of the conditioning vector. The embedders $\{\mathcal{E}_{i}\}_{i=1}^{K}$ are trained jointly with the denoising network end-to-end; they are not independently pretrained feature extractors. As a result, we can selectively null-tokenize or modulate individual attributes at inference time, enabling fine-grained control. Throughout the rest of the paper, we denote the semantic attribute vector as $\mathbf{pa}$ and the corresponding embedding vector as $\mathbf{c}$, as defined in Eq. 10.
3.2 Group-wise Formulation of FCFG
To overcome the limitations of CFG and enable more precise, causally aligned control in counterfactual image generation, we propose Factored Classifier-Free Guidance (FCFG). Let $\mathbf{pa}{=}(pa_{1},\dots,pa_{K})$ denote the vector of semantic parent attributes. Rather than applying a single guidance weight uniformly to the entire conditioning vector, we partition semantic attributes $\mathbf{pa}$ into $M$ disjoint groups $\mathbf{pa}^{(1)},\mathbf{pa}^{(2)},\dots,\mathbf{pa}^{(M)}$, and apply a separate guidance weight $\omega_{m}$ to each group.
Proposition 3.1 (Proxy Posterior for FCFG).
Under the assumption that the groups $\mathbf{pa}^{(1)},\dots,\mathbf{pa}^{(M)}$ are conditionally independent given the latent variable $\mathbf{x}_{t}$, for any time $t$, that is: $p(\mathbf{pa}{\mid}\mathbf{x}_{t}){=}\prod_{m=1}^{M}p(\mathbf{pa}^{(m)}{\mid}\mathbf{x}_{t})$, we obtain the following factorized proxy posterior:
$p^{\omega}(\mathbf{x}_{t}\mid\mathbf{pa})\propto p(\mathbf{x}_{t})\prod_{m=1}^{M}p(\mathbf{pa}^{(m)}\mid\mathbf{x}_{t})^{\omega_{m}},$ (11)
where $\omega_{m}{\geq}0$ controls the guidance strength for each group.
A complete derivation and score-based justification for this proxy posterior is provided in App. B. The corresponding guided update used in score-based diffusion sampling is then given by:
$\nabla_{\mathbf{x}_{t}}\log p^{\omega}(\mathbf{x}_{t}{\mid}\mathbf{pa}){=}\nabla\log p(\mathbf{x}_{t})$ (12)
${+}\sum_{m=1}^{M}\omega_{m}\big(\nabla\log p(\mathbf{x}_{t}{\mid}\mathbf{pa}^{(m)}){-}\nabla\log p(\mathbf{x}_{t})\big).$
[…]
The final guided score used in the diffusion model is computed as follows:
$\boldsymbol{\epsilon}_{\mathrm{FCFG}}$ $(\mathbf{x}_{t},t,\mathbf{c})=\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\varnothing)$ (14)
[…]
Figure 1: Comparison of $\Delta$ metrics under different interventions in CelebA-HQ. Left: Intervention on Smiling. Right: Intervention on Young. Both use baseline $\omega{=}1.0$. Under global CFG, increasing $\omega$ boosts the intended attribute but amplifies non-target ones. FCFG achieves similar improvements on the target attribute while mitigating amplification. See App. F.2 for full results.
We emphasize that the model is trained with standard classifier-free dropout, where the entire conditioning vector is replaced by the null token, rather than with all possible partially masked conditioning patterns. Thus, FCFG introduces a mild train–test mismatch at inference time by evaluating the denoiser on group-masked embeddings. Empirically, we do not observe systematic artifacts or instability from this mismatch, but it may become more pronounced with many groups, stronger guidance weights, or more severe masking patterns.
Figure 2: Counterfactual generations in CelebA-HQ ($64{\times}64$). Each row compares global CFG (left) and FCFG (right) across guidance weights. Top: under do(Male=no), global CFG also amplifies Smiling, although Smiling is not intervened upon. Middle: under do(Young=no), global CFG produces increasingly strong age-related changes while also altering facial expression and identity beyond the intended age intervention. Bottom: under do(Smiling=yes), global CFG makes the subject appear older, adds glasses, and alters identity. FCFG mitigates these unintended changes and better preserves invariant attributes. See App. F.3 for more visual results.
@@ 3.2 Group-wise Formulation of FCFG
The proposed FCFG framework is highly flexible, as it allows arbitrary groupings of attributes, regardless of whether attributes within a group are mutually independent or not. Consequently, strongly entangled attributes should be placed in the same group rather than separated into different guidance groups. For the factorized proxy posterior in Proposition 3.1, the required assumption is conditional independence across groups given the latent variable $\mathbf{x}_{t}$. This flexibility enables a wide range of configurations. For instance, setting M=1M{=}1 recovers standard global CFG, while increasing $M$ provides progressively finer-grained control, including per-attribute guidance (M=KM{=}K) as an extreme case where we assume all attributes are independent of each other. Algorithm 4 summarizes the general group-wise FCFG counterfactual prediction procedure.
3.3 FCFG for Counterfactual Generation
@@
We now detail how FCFG is straightforwardly integrated into DDIM-based counterfactual inference.
@@
Abduction. The abduction step proceeds as in Eq. 4, where the conditioning vector $\mathbf{c}$ is obtained by embedding the semantic parent attributes $\mathbf{pa}$ using the attribute-split embedder defined in Eq. 10.
@@
Prediction. The prediction step uses the FCFG-guided reverse DDIM trajectory:
[…]
For counterfactual generation, we adopt a two-group partitioning of attributes based on the assumed causal graph. The affected group $\mathbf{pa}^{\text{aff}}$ contains attributes directly intervened upon and their descendants, while the invariant group $\mathbf{pa}^{\text{inv}}$ comprises attributes expected to remain unchanged. For the proxy-posterior interpretation in Proposition 3.1, these groups are assumed conditionally independent given the latent $\mathbf{x}_{t}$. In practice, this grouping is informed by the assumed causal graph: attributes directly intervened upon and their descendants are assigned to the affected group, while attributes outside the intervention’s causal effect are assigned to the invariant group. Under this setup, Eq. 14 uses M=2M{=}2 groups with separate guidance weights $\omega_{\text{aff}}$ and $\omega_{\text{inv}}$. The two-group counterfactual prediction procedure is summarized in Algorithm 5.
Note that the two-group partition is only one possible choice: guidance can also be separated at a finer-grained level, including the attribute level, provided conditional independence holds across groups. We present such extensions, including multi-attribute interventions and attribute-wise configurations, in App. I, which further demonstrate the generality and flexibility of FCFG. The proposed factoring framework can be integrated with recent CFG variants such as CFG++ and APG by applying the group-wise decomposition to their guided score formulations (see Apps. J and K).
@@ 4 Experiments
In this section, we demonstrate the benefits of the proposed approach across three public datasets. For each dataset, we train a diffusion model with the same architecture and training protocol, detailed in App. D. We first compare FCFG against the standard CFG baseline, and then further contextualize it against existing counterfactual generation methods, including HVAE, HVAE-soft, and SA-DCG. Unless otherwise specified, settings labeled as $\omega{=}X$ correspond to standard CFG with a global guidance weight. In contrast, configurations denoted by $\omega_{\text{aff}}{=}X$, $\omega_{\text{inv}}{=}Y$ represent the two-group FCFG, where separate guidance weights are applied to the intervened and invariant attribute groups, respectively.
Following Monteiro et al. (2023); Melistas et al. (2024), we primarily evaluate counterfactual quality using two metrics. Effectiveness ($\Delta$): Measured by a pretrained classifier as the change in AUROC for target or invariant attributes relative to $\omega=1.0$ (no CFG). Higher $\Delta$ on the intervened attribute indicates stronger target-attribute change, while large $\Delta$ on invariant attributes indicates unintended amplification. Reversibility: Assesses how well counterfactuals can be reversed to the original image using inverse interventions. We report MAE and LPIPS; lower values indicate better identity preservation. See App. A.2 for details.
4.1 Main Results
@@
Figure 3: Reversibility analysis in CelebA-HQ $(64{\times}64)$. Left: Quantitative evaluation of how well the original image is recovered after generating a counterfactual and mapping it back to the original condition under do(Smiling). Right: A qualitative example showing a counterfactual generated under do(Male) and its reconstruction after reversing the intervention with CFG and our FCFG.
CelebA-HQ. We begin our empirical evaluation of FCFG on the CelebA-HQ dataset (Karras et al., 2017), using Smiling, Male, and Young as independent binary attributes. We adopt this simplified setup to isolate inference-time failures of standard CFG in a controlled setting. Under this designed scenario, unintended changes in non-intervened attributes can be attributed to attribute amplification rather than valid causal effects. Refer to App. F.1 for more dataset details.
Fig. 1 presents the $\Delta$ metrics under different guidance strategies for two separate interventions, namely do(Smiling) and do(Young). As the global guidance weight $\omega$ of CFG increases (left to right side of each plot), the $\Delta$ of the intervened attribute improves, but so do the $\Delta$ values of attributes that should remain invariant, indicating undesirable spurious amplification. In contrast, the right side of each plot shows results for FCFG, where distinct weights are applied to affected ($\omega_{\text{aff}}$) and invariant ($\omega_{\text{inv}}$) attribute groups. This decoupled formulation achieves comparable or stronger improvement on the intervened attribute while keeping the others stable, validating the ability of FCFG to produce more disentangled and effective counterfactuals purely at inference time.
Fig. 2 illustrates how global CFG can introduce unintended changes by uniformly amplifying all conditioning signals, even when only one attribute is meant to change. In the top row, applying do(Male=no) with increasing $\omega$ inadvertently amplifies Smiling; in the middle row, do(Young=no) changes non-target facial attributes and expression in addition to age; and in the bottom row, do(Smiling=yes) introduces changes to age, identity, and even adds glasses. These unintended shifts stem from global CFG treating all attributes equally. In contrast, FCFG applies decoupled guidance across attribute groups, assigning stronger weights to attributes affected by the intervention while preserving invariant factors. This results in counterfactuals that more faithfully reflect the intended change while better preserving identity and consistency in invariant attributes.
Fig. 3 evaluates the reversibility of counterfactuals in CelebA-HQ. The left panel shows MAE and LPIPS when recovering the original after applying and then reversing an intervention (e.g., do(Smiling)). With global CFG, errors grow as guidance strength increases, while FCFG yields consistently lower values for the same settings, improving recovery. The right panel shows a qualitative example under do(Male), where global CFG amplifies non-intervened attributes (e.g., Young), making the reversed image appear older. In contrast, FCFG applies strong guidance only to intervened attributes, mitigating amplification and producing more faithful, disentangled, and reversible counterfactuals. More reversibility results are provided in Apps. F.2 and F.3.
@@
Figure 4: Evaluation of counterfactual generation on EMBED ($192{\times}192$). Left: $\Delta$ metrics showing the effect of do(circle). FCFG improves target intervention effectiveness while suppressing spurious shifts in non-intervened attributes. Right: A visual example showing the input image, the counterfactual under do(density), the reversed image, and their difference maps (CF/Rev.-input). See App. G.2 for full quantitative results and App. G.3 for more visual results.
EMEBD. We evaluate FCFG on the EMBED (Jeong et al., 2023) breast mammography dataset. For our experiments, we define a binary circle attribute based on the presence of circular skin markers, and a binary breast density label, where categories A and B are grouped as low and categories C and D as high. See App. G.1 for dataset details.
Fig. 4 presents results for counterfactual generation on EMBED. The left bar plot reports $\Delta$ effectiveness metrics, measuring how classifier performance changes relative to the baseline. While global CFG improves effectiveness for the target attribute (circle), it also increases effectiveness on non-intervened attributes such as density, indicating unintended attribute amplfication. FCFG mitigates this by applying selective guidance, maintaining stable performance on non-target attributes. The figure on the right illustrates a key example: applying do(density) under global CFG unintentionally amplifies the presence of circular skin markers, as evidenced by the increased number of visible circles in both the counterfactual and reversed images. This is suppressed under FCFG, where circle features remain unchanged in both counterfactual and reversed images.
@@
Figure 5: Evaluation of counterfactual generation on MIMIC ($192{\times}192$). Left: $\Delta$ metrics showing the effect of do(finding). FCFG improves target intervention effectiveness while suppressing spurious shifts in non-intervened attributes. Right: A visual example showing the input image, the counterfactual under do(finding), the reversed image, and their difference maps (CF/Rev.-input). See App. H.2 for full quantitative results and App. H.3 for more qualitative results.
MIMIC. We evaluate our method on the MIMIC-CXR dataset (Johnson et al., 2019). We follow the dataset splits and filtering protocols from Ribeiro et al. (2023), and focus on the binary disease label of pleural effusion. The underlying causal graph in Ribeiro et al. (2023) includes four attributes: race, sex, finding, and age. We adopt this setup, but since our goal is to study attribute amplification caused by CFG, we focus on sex, race, and finding, which we assume to be mutually independent for the purposes of our analysis. Refer to App. H.1 for dataset details.
Fig. 5 presents an evaluation of counterfactual generation in MIMIC-CXR, highlighting the advantages of our proposed FCFG. The bar plot on the left shows $\Delta$ metrics that quantify the change in effectiveness relative to the baseline $\omega{=}1.0$. While global CFG improves effectiveness for the intervened variable (finding) as expected, it also introduces substantial spurious shifts in non-intervened attributes such as race and sex, revealing unwanted attribute amplification. In contrast, FCFG achieves comparable or higher intervened effectiveness while suppressing spurious amplification, demonstrating more precise and controlled generation. On the right, we show a qualitative example of a counterfactual generated under do(finding), its reversed reconstruction, and their corresponding difference maps. Compared to global CFG, our method yields localized, clinically meaningful changes in counterfactuals and better identity preservation in the reversed image.
4.2 Comparison with Existing Counterfactual Methods
To further contextualize FCFG against existing counterfactual generation methods, we compare with HVAE (Ribeiro et al., 2023), HVAE-soft (Xia et al., 2024), and SA-DCG (Rasal et al., 2025); detailed numerical results are provided in App. E.1. These methods also rely on structured causal information, but incorporate it differently: HVAE and HVAE-soft use predictor-based counterfactual training objectives, while SA-DCG uses a diffusion-autoencoder framework with semantic abduction. In contrast, FCFG modifies only the inference-time guidance rule of a trained conditional diffusion model.
Figure 6: Qualitative results for do(Smiling, Male, Young). Since all three attributes are intervened, no invariant attributes remain; therefore, global CFG and the two-group variant of FCFG are equivalent, as a single guidance weight is applied to all attributes. Attribute-wise FCFG further provides flexibility by enabling independent control over each attribute; setting ${\omega_{\text{s}}}{=}{\omega_{\text{m}}}{=}{\omega_{\text{y}}}{=}2.5$ recovers the global/two-group configuration. See App. I.2 for more quantitative and qualitative results.
Across CelebA-HQ and MIMIC-CXR, FCFG achieves a stronger target/off-target trade-off: it maintains or improves intervention effectiveness on the target attribute while substantially reducing off-target amplification. On CelebA-HQ under do(Smiling), SA-DCG improves reversibility relative to standard CFG, but still exhibits off-target amplification under global guidance. At the stronger guidance setting, SA-DCG reaches a target AUC change of +12.9+12.9 but also increases the off-target Male AUC by +3.0+3.0, whereas FCFG reaches a slightly higher target change of +13.1+13.1 while reducing the off-target change to 1.5-1.5, with better reversibility measured by MAE and LPIPS.
On MIMIC-CXR, FCFG also achieves comparable or stronger target intervention than HVAE, HVAE-soft, and CFG, while reducing off-target change by a large margin. For example, under do(sex), FCFG achieves a target AUC change of +7.6+7.6 with only +0.6+0.6 off-target change, compared with +9.0+9.0, +5.7+5.7, and +10.7+10.7 off-target change for HVAE, HVAE-soft, and CFG, respectively. Under do(finding), FCFG achieves the strongest target change of +18.8+18.8 while reducing off-target change to +0.6+0.6. These results support our central claim that the global guidance mechanism itself is an important source of attribute amplification, and that factorizing guidance provides a simple inference-time remedy.
We further evaluate general image quality using FID and analyze the practical cost of FCFG. On CelebA-HQ, FCFG consistently improves FID over global CFG across do(Smiling), do(Male), and do(Young), suggesting that reducing off-target amplification also helps preserve image realism. FCFG also offers a favourable simplicity–control trade-off: unlike HVAE and HVAE-soft, it does not require an additional predictor-based counterfactual training objective, and unlike SA-DCG, it avoids a heavier diffusion-autoencoder framework and identity-preserving inference-time optimization. Full baseline comparisons, FID results, dropout ablations, and practical-cost comparisons are provided in App. E.
4.3 Generality and Compatibility of FCFG
Multi-attribute intervention. To illustrate that FCFG extends beyond the two-group setting used in our main experiments, we consider a three-attribute intervention do(Smiling, Male, Young). In this setting, all attributes are directly intervened upon, and no invariant attributes remain; as a result, the standard two-group formulation collapses to global guidance, since a single guidance weight is applied to all attributes. As shown in Fig. 6, attribute-wise FCFG instead assigns independent guidance weights ($\omega_{\text{s}},\omega_{\text{m}},\omega_{\text{y}}$) to each attribute. While symmetric weights recover the global configuration, asymmetric settings demonstrate the additional flexibility of FCFG in selectively controlling individual semantic factors. Further discussion and results on multi-attribute interventions are provided in App. I.
Compatibility with advanced guidance schemes. FCFG is compatible with advanced guidance variants such as CFG++ (Chung et al., 2025) and APG (Sadat et al., 2025). While both methods still suffer from attribute amplification in counterfactual generation, integrating FCFG mitigates this issue. Algorithmic descriptions and quantitative and qualitative results are provided in Apps. J and K.